OpenStudy (anonymous):

An electric vehicle starts from rest and accelerates at a rate of 2.6 m/s2 in a straight line until it reaches a speed of 10 m/s. The vehicle then slows at a constant rate of 1.0 m/s2 until it stops. (a) How much time elapses from start to stop?(b) How far does the vehicle travel from start to stop?

OpenStudy (roadjester):

\[\large a_1=2.6 m/s^2\] \[\large a_2=-1 m/s^2\] \[\large v_i=0 m/s\] \[\large v_f=10 m/s\] this is mostly a substitution problem using your kinematic formulas

OpenStudy (roadjester):

\[\large x_f = x_i + v_i t + \frac 1 2 a t^2\]

OpenStudy (anonymous):

what is x, if v is velocity, t is time and a is acceleration?

OpenStudy (roadjester):

x is position instead of p p is used to represent linear momentum which you will probably learn about later on

OpenStudy (anonymous):

so what numbers or constants can i plug in for x?

OpenStudy (roadjester):

ok, so you start at rest correct?

OpenStudy (roadjester):

what values can you imply if you start from rest?

OpenStudy (roadjester):

and which are most convenient?

OpenStudy (anonymous):

yes which equals 0 , right?

OpenStudy (roadjester):

remember, whenever possible, you want to use the most convenient values. this does not necessarily imply 0, but typically does because 0 allows you to cancel out terms

OpenStudy (anonymous):

so then what would the value be at rest if it isnt zero because from understanding rest means not moving which would equal 0

OpenStudy (roadjester):

okay, I was too vague there

OpenStudy (roadjester):

when you are at rest, VELOCITY IS ALWAYS 0 position however is a matter of convenience. typically, if you use 0 that will cancel out the x-initial term, but sometimes maybe you are using a different scale or whatnot does that make sense?

OpenStudy (anonymous):

okay i think i understand but then what kind of factors determine position?

OpenStudy (roadjester):

ok, do you know the four basic kinematic formulas?

OpenStudy (anonymous):

no

OpenStudy (roadjester):

velocity as a function of time \[\large v_f = v_i + at\] position as a function of velocity and time \[\large x_f=x_i + \frac 1 2 (v_i+v_f)t\] position as a function of time \[\large x_f = x_i + v_i t+ \frac 1 2 a t^2\] velocity as a function of position \[\large v_f^2 = v_i^2 + 2a(x_f-x_i)\]

OpenStudy (roadjester):

the important thing to remember though is that all of these equations assume one-dimensional motion

OpenStudy (roadjester):

you'll probably get into two dimensional motion later on

OpenStudy (roadjester):

as of right now, I would recommend the first equation, since you are looking for time

OpenStudy (roadjester):

but remember you need to do it twice, once for acceleration, and once for deacceleration and add the two times

OpenStudy (roadjester):

from there, you could probably apply any of the first three to get distance travelled

OpenStudy (anonymous):

what do you mean by "and add the two times"

OpenStudy (anonymous):

do you mean add the sums together?

OpenStudy (roadjester):



OpenStudy (roadjester):

you need to find t1 which is when you are accelerating or speeding up then once you hit v=10m/s, you're going to deaccelerate. But remember than when you deacclerate, v_i and v_f switch

OpenStudy (roadjester):

does it make sense? do you understand what you need to do?

OpenStudy (anonymous):

okay so i know for time 1 the answer is 10/2.6 seconds but how do i find time 2

OpenStudy (roadjester):

ok so now, you're going at v=10 m/s so that is now your initial velocity you're going to keep going until you stop so final velocity is 0 that's why i said you swap the velocities your accleration is now -1m/s^2

OpenStudy (anonymous):

okay so then the time 2 is 10 s

OpenStudy (roadjester):

is it correct?

OpenStudy (anonymous):

do you mean the second time in seconds?

OpenStudy (roadjester):

no. I don't have the answer. I was wondering if you had the correct answer

OpenStudy (anonymous):

\[x _{f}=x _{i}+0+(1/2)(2.6)(10/2.6)^2\]

OpenStudy (anonymous):

i still dont know what x is ? i understand its a position but what is its value?

OpenStudy (roadjester):

ok, so let's say for a minute that the total distance traveled is 10 units (on an x-y axis) does it matter if you start from -5 and go to 5, or 0 and go to 10? or even start at 52 and go to 62?

OpenStudy (anonymous):

no it doesnt matter, 10 units is 10 units

OpenStudy (roadjester):

\[\Delta x= x_f-x_i\] delta x is also distance travelled so if you don't care where you start, if you move x_i to the left of the equal sign, do you agree that you get the distance traveled?

OpenStudy (anonymous):

do you mean \[\Delta(x)-x _{i}=x _{f}\]

OpenStudy (roadjester):

\[\large x_f=x_i +v_i t + \frac 1 2 a t^2\] \[\large x_f - x_i =x_i -x_i+v_i t + \frac 1 2 a t^2\] \[\large \Delta x=v_i t + \frac 1 2 a t^2\]

OpenStudy (roadjester):

this make sense?

OpenStudy (anonymous):

yes , so now i have to find what delta x is

OpenStudy (roadjester):

no you don't, you already found delta x

OpenStudy (roadjester):

look at your earlier post

OpenStudy (anonymous):

delta x is 10?

OpenStudy (roadjester):

actually i used 10 as an example, I meant YOUR post, not mine

OpenStudy (roadjester):

the one where you substituted numbers

OpenStudy (anonymous):

okay so then delta x =19.23, is that my answer?

OpenStudy (roadjester):

first of all was your total time 13.8s? assuming my approach was correct, that was my answer

OpenStudy (anonymous):

yes my total time was 13.8. was i supposed to use the total time to complete the equation because i only used the time for acceleration at 2.6 m/s

OpenStudy (roadjester):

I was just checking but ok yes delta x 1 is 19.2 m

OpenStudy (roadjester):

now think about this for a second, earlier you said that 10 units is 10 units no matter where you start. so which location is the most convenient initial position such that the final position IS delta x?

OpenStudy (anonymous):

initial position could be zero

OpenStudy (roadjester):

remember, in kinematics, these equations are a matter of convenience, you can start where you want but some locationns will make your life hell

OpenStudy (roadjester):

YES!

OpenStudy (anonymous):

so then 19.2 =\[x _{f}-0\]

OpenStudy (anonymous):

making \[x _{f}=0\]

OpenStudy (roadjester):

correction x_i=0, not x_f

OpenStudy (roadjester):

also remember that you can only do this when you start at rest

OpenStudy (anonymous):

i mean \[x _{f}=19.2\]

OpenStudy (roadjester):

so for the second part when you deaccelerate your initial position is not 0, it's 19.2

OpenStudy (roadjester):

make sense?

OpenStudy (anonymous):

yes because you are leaving from the postiion you went to

OpenStudy (roadjester):

just as a side note, in THIS senario, there is no difference, but you won't always be that lucky so if you make the mistake of assuming that your initial velocity is 0 and proceed, you may get lucky, but sooner or later, a senario will come up in which it comes back to bite you

OpenStudy (roadjester):

over time you become more experienced and can tell when it doesn't matter

OpenStudy (anonymous):

so how do i know what the initital velocity is if its not zero?

OpenStudy (roadjester):

it's 10

OpenStudy (roadjester):

it's the velocity you were at when you started to deaccelerate

OpenStudy (anonymous):

oh okay so do i have to use the formula \[v _{f}=v _{i}+at\]

OpenStudy (roadjester):

remember to focus on what you're looking for now you're looking for distance 2

OpenStudy (anonymous):

which is delta x, right?

OpenStudy (roadjester):

also remember that you are looking for x_final x-initial is 19.2 so you can just leave x_final on the left and x-initial on the right

OpenStudy (anonymous):

\[x _{f}=19.2+(10)(13.8)(1/2)a(13.8)^2\]

OpenStudy (roadjester):

let's call your starting point A let's call B the point where your velocity hit 10m and your acceleration changes to -1m/s^2 lets call C the point you finally stop Time time it took to go from point B to point C was 10 s whereas 13.8 was the time it too to go from A to C based on this, which time should you use? remember, your initial position right now is 19.2, which is where you stopped when going from A to B and where B starts

OpenStudy (roadjester):

for going from B to C

OpenStudy (anonymous):

so then the time is 10 s and the acceleration is -1 m/s

OpenStudy (roadjester):

yes

OpenStudy (anonymous):

then the equation is \[x _{f}=19.2+(10)(10)+(1/2)(-1)(10)^2\]

OpenStudy (roadjester):

that's what i got

OpenStudy (anonymous):

the answer is 69.2?

OpenStudy (roadjester):

well that's what I got one sec @LastDayWork could you do me a favor and check whatI did? relativity's burned me out so maybe you'll catch something I did wrong

OpenStudy (anonymous):

you are right, i checked on my homework . thank you so much !!!!!

OpenStudy (roadjester):

well okay then and you're welcome @LastDayWork thanks

OpenStudy (lastdaywork):

:)